# Evaluating Multi-Hop Reasoning in Large Language Models: A Chemistry-Centric Case Study

### v1

@@ Abstract
In this study, we introduced a new benchmark consisting of a curated dataset and a defined evaluation process to assess the compositional reasoning capabilities of large language models within the chemistry domain. We designed and validated a fully automated pipeline, verified by subject matter experts, to facilitate this task. Our approach integrates OpenAI reasoning models with named entity recognition (NER) systems to extract chemical entities from recent literature, which are then augmented with external knowledge bases to form a comprehensive knowledge graph. By generating multi-hop questions across these graphs, we assess LLM performance in both context-augmented and non-context augmented settings. Our experiments reveal that even state-of-the-art models face significant challenges in multi-hop compositional reasoning. The results reflect the importance of augmenting LLMs with document retrieval, which can have a substantial impact on improving their performance. However, even perfect retrieval accuracy with full context does not eliminate reasoning errors, underscoring the complexity of compositional reasoning. This work not only benchmarks and highlights the limitations of current LLMs but also presents a novel data generation pipeline capable of producing challenging reasoning datasets across various domains. Overall, this research advances our understanding of reasoning in computational linguistics.
@@ 1 Introduction
    We contribute a challenging benchmark for Multi-hop QA in chemistry
The complete Q&A dataset, together with the evaluation code, is available  [^1] https://github.com/MohammadKhodadad/ChemKGMultiHopQA.
@@ 2 Background and Related Works
Multi-hop question answering (QA) has evolved as a key method to evaluate the multi-step reasoning abilities of large language models [25]. Answering multi-hop questions requires integrating multiple pieces of evidence. Traditionally, datasets such as HotpotQA [19], WikiHop, and MedHop [26] were manually or semi-automatically curated through crowd-sourcing and knowledge-base relations. While these approaches yield high-quality, human-validated questions, they are resource-intensive.
More recently, LLMs have been leveraged to generate multi-hop datasets automatically. In many cases, single-hop QA pairs are first generated and later merged using entity linking techniques, a common approach that connects individual entities across questions. For example, the MuSiQue framework [27] fuses two QA pairs by linking a named entity from the first answer to the subsequent question, thereby forming a chain of reasoning. Other methods, such as MultiHop-RAG [28], extend this paradigm by incorporating retrieval-augmented generation (RAG) to paraphrase factual sentences and group them based on shared topics, reflecting the diverse strategies that are emerging in multi-hop QA generation.
@@
The chemical sciences pose distinct challenges for multi-hop QA due to the need for expert domain knowledge. Only a small fraction of HotpotQA’s questions (roughly 900) are chemistry-focused, limiting both domain relevance and topical currency. Furthermore, they tend to be limited to 2-hop reasoning which limits the difficulty level of the questions. More specialized datasets, such as ChemLitQA [21], provide around 1,000 single-hop and 700 multi-hop question-answer pairs generated using LLMs based on ChemRxiv papers. Although the ChemLitQA-multi dataset includes multi-hop questions, these are typically defined by a single linked entity across all hops. This limitation has led the authors of [21] to emphasize the need for future studies focused on developing more challenging multi-hop QA benchmarks in the field.
Similarly, the chemistry subset of the OlympicArena benchmark [22] provides high-level Olympiad problems that are few in number, and are not explicitly designed for multi-hop reasoning.
@@ 3 Methodology
Our methodology consists of three main components: knowledge graph generation, multi-hop question-answer generation, and evaluating state-of-the-art large language models on the question-answering task. The first two components are described in detail in this section, and the last one is explained in the next section.
@@ 3.1 Knowledge Graph Generation
Next, we applied named entity recognition (NER) models to these text chunks to identify chemical entities. In particular, we utilized an NER model [34] that leverages a PubMedBERT architecture [35] fine-tuned on various chemical datasets. To ensure that the extracted entities were specific, verified, and chemically relevant, we utilized OpenAI’s gpt-4o to review and refine the outputs. The same model was also employed to extract relations between these verified entities, forming triplets that capture the interactions and associations present in the text. Additionally, large language models were utilized to extract descriptive features from textual data associated with each entity. To enrich the nodes further during the construction of the knowledge graph, supplementary information from Wikipedia and the PubChem dataset [36] was integrated. Consequently, the finalized knowledge graph comprises nodes representing chemical entities, enhanced by metadata and descriptive annotations from these sources, as well as edges representing the relationships extracted from the textual data. Figure 1 illustrates the procedure followed to generate the knowledge graph.
@@ 3.2 Multi-hop Question-Answer Generation
These individual questions were then combined into a single multi-hop question using OpenAI’s o3-mini model. Importantly, the final aggregated question was constructed to begin with the last sub-question and chain the entities up to the $\text{entity}_{1}$ of the first relation, ensuring that the final answer corresponds to the answer of the first question. During the verification phase, each one-hop question was first reviewed for clarity, relevance to chemistry, and alignment with the corresponding text that provided the answer. The multi-hop question was then assessed through an additional evaluation step, ensuring that its logical flow effectively led to the final answer. An LLM-based verification process was employed to confirm factual accuracy, answerability based on available context and metadata, and the logical coherence of the sub-questions. Feedback from domain experts was continuously incorporated into the prompts to enhance verification accuracy. To minimize ambiguity, questions that were answered incorrectly by all evaluated models were excluded from the benchmark. Figure 2 illustrates the detailed pipeline of Multi-hop QA generation.
4 Experiments and Results
To access the selected models for evaluation in this experiment, we used different API providers: (i) all tested OpenAI models (gpt-4o, gpt-4o-mini, o1-mini and o3-mini) are accessed via the OpenAI platform; (ii) Amazon Bedrock Platform has been used to access Anthropic Sonnet 3.7 (with and without extended thinking), Anthropic Sonnet 3.5 V2, Mistral Large, DeepSeek R1 and Llama 3.3 70B Instruct; and (iii) Google Gemma 3 27B, Qwen QwQ 32B, and DeepSeek R1 Distill Qwen 32B are accessed via the OpenRouter Platform and operate at bf16 precision. All of these models can perform function-calling tool use, so they were instructed to produce valid JSON outputs to ensure consistency and enable automated validation. All models are evaluated in two settings: with and without provided context. The first scenario reflects performance when the models are paired with an ideal retrieval-augmented generation (RAG) system, while the second scenario relies on the model’s internal memory to answer the questions. After parsing the JSON, we checked whether the output was an exact match to ground truth. If not, OpenAI GPT-4o was instructed to perform a binary assessment; determining whether the answer was correct or not, to calculate the Correctness Rate (%) metric. Our dataset comprises 971 questions spanning 1 to 4 hops (On average 245 questions per hop), generated using the approach described in Section 3.2.
@@ 4.1 Models Performance
Figure 3 illustrates the performance of 13 large language models evaluated with respect to correctness rate, cost, and latency in both context-provided and context-not-provided setup. In our performance evaluation, the Llama 3.3 70B Instruct and GPT-4o models achieved the lowest cost and demonstrated notably low latency, but they also registered the lowest correctness rate, making them cost-efficient yet less accurate options. In contrast, Claude Sonnet 3.7 (with extended thinking) achieved the highest correctness rate, albeit at the expense of significantly higher cost and latency. Meanwhile, both Qwen QWEN 32B and Deepseek R1 Distil QWEN 32B maintained a favorable balance between cost and correctness rate when the context is provided – i.e. equipped with a perfect RAG system –, though they incurred above-average latency. Claude Sonnet 3.7 was found to have the highest correctness rate in non-reasoning models.
@@
Figure 3: Performance of selected models based on correctness rate, cost, and latency. The cost axis is log-scaled to emphasize differences. The y-axis represents the percentage of questions answered correctly by each model, while the size of the dots indicates the average latency for each model when answering questions. The top panel displays results for setups where context is provided to the models, while the bottom panel presents results for setups without context. Note that the horizontal axis in both panels has the same range, while the vertical axes ranges differ.
@@
In Context not provided setup, open-source reasoning models (R1 Distill Qwen, QWQ-32B, and R1) tend to have lower performance ranking compared to other models. In contrast, for OpenAI models we observed an opposite trend, which may indicate potentially richer pre-training data. For Claude 3.7, using extended thinking did not lead to better performance compared to the no-thinking setup when the context is not provided to the model; instead, it increased the output tokens and cost. For a comprehensive breakdown of model performance and experimental settings, refer to Table 1, which details metrics such as correctness rate, latency, and token usage. Note that the output details for reasoning models also include the reasoning tokens.
4.2 Comparison with HotpotQA
@@
To demonstrate that large language models, even those designed for reasoning, often struggle with domain-specific multi-hop questions, we created a chemistry-related subset of HotpotQA [19], a well-known general text benchmark primarily sourced from Wikipedia. We sampled chemistry questions by starting from Wikipedia’s Chemistry category, recursively exploring its subcategories (up to three levels), and filtering HotpotQA based on exact title matches. To maintain consistency with our evaluation scheme, we excluded distractors and included only supporting documents as context. Figure 4, illustrates the average performance of all models on each dataset under two conditions: with context provided and without context in the prompt. The results indicate that when context is provided, models achieve similar performance, with our benchmark resulting in marginally lower average performance and reduced variability among the models’ output. In the setup without context, the models found our benchmark more challenging compared to the HotpotQA chemistry subset. This observation may be due to the fact that HotpotQA was exclusively built from Wikipedia, which has been utilized in the pre-training of all evaluated models, while the new benchmark is constructed from more recent ChemRxiv papers enriched with PubChem and Wikipedia.
@@
Figure 4: Comparison of LLMs performance on the chemical subset of HotPotQA with the curated QA dataset from this study. Error bars reflect the standard error of the mean (S.E.M) over evaluated models.
@@ 5 Analysis and Ablation
This section presents detailed ablation studies and analyses of the benchmark. First, we examine how context availability and test-time reasoning affect model performance and efficiency. Next, we explore how the number of reasoning hops, an indicator of question difficulty, impacts model correctness and the number of tokens needed to generate an answer.

### v2

@@ Abstract
In this study, we introduced a new benchmark consisting of a curated dataset and a defined evaluation process to assess the compositional reasoning capabilities of large language models within the chemistry domain. We designed and validated a fully automated pipeline, verified by subject matter experts, to facilitate this task. Our approach integrates OpenAI reasoning models with named entity recognition (NER) systems to extract chemical entities from recent literature, which are then augmented with external knowledge bases to form a comprehensive knowledge graph. By generating multi-hop questions across these graphs, we assess LLM performance in both context-augmented and non-context augmented settings. Our experiments reveal that even state-of-the-art models face significant challenges in multi-hop compositional reasoning. The results reflect the importance of augmenting LLMs with document retrieval, which can have a substantial impact on improving their performance. However, even perfect retrieval accuracy with full context does not eliminate reasoning errors, underscoring the complexity of compositional reasoning. This work not only benchmarks and highlights the limitations of current LLMs, but also presents a novel data generation pipeline capable of producing challenging reasoning datasets across various domains. Overall, this research advances our understanding of reasoning in computational linguistics.
@@ 1 Introduction
    We contribute a challenging benchmark for Multi-hop QA in chemistry, reviewed by our domain experts.
@@ 2 Background and related works
Multi-hop Question Answering.
Multi-hop question answering (QA) has evolved as a key method to evaluate the multi-step reasoning abilities of large language models [25]. Answering multi-hop questions requires integrating multiple pieces of evidence. Traditionally, datasets such as HotpotQA [19], WikiHop, and MedHop [26] were manually or semi-automatically curated through crowd-sourcing and knowledge-base relations. While these approaches yield high-quality, human-validated questions, they are resource-intensive. More recently, LLMs are leveraged to automatically generate multi-hop datasets. In many cases, single-hop QA pairs are first generated and later merged using entity linking techniques, a common approach that connects individual entities across questions. For example, the MuSiQue framework [27] fuses two QA pairs by linking a named entity from the first answer to the subsequent question, thereby forming a chain of reasoning. Other methods, such as MultiHop-RAG [28], extend this paradigm by incorporating retrieval-augmented generation (RAG) to paraphrase factual sentences and group them based on shared topics, reflecting the diverse strategies that are emerging in multi-hop QA generation.
@@
The chemical sciences pose distinct challenges for multi-hop QA due to the need for expert domain knowledge. Only a small fraction of HotpotQA’s questions (roughly 900) are chemistry-focused, limiting both domain relevance and topical currency. Furthermore, they tend to be limited to 2-hop reasoning, which limits the difficulty level of the questions. More specialized datasets, such as ChemLitQA [21], provide around 1,000 single-hop and 700 multi-hop question-answer pairs generated using LLMs based on ChemRxiv papers. Although the ChemLitQA-multi dataset includes multi-hop questions, these are typically defined by a single linked entity across all hops. This limitation has led the authors of [21] to emphasize the need for future studies focused on developing more challenging multi-hop QA benchmarks in the field.
Similarly, the chemistry subset of the OlympicArena benchmark [22] provides high-level Olympiad problems that are few in number, and are not explicitly designed for multi-hop reasoning. As another recent effort, [29] introduces ChemBench, an automated benchmark of over 2,700 expert-validated chemistry QA pairs to systematically assess LLMs’ chemical reasoning against human experts.
@@ 3 Methodology
Our methodology consists of three main components: knowledge graph generation, multi-hop question-answer generation, and evaluating state-of-the-art large language models on the question-answering task. The first two components are described in this section, and the last one is explained in the next section.
@@ 3.1 Knowledge Graph Generation
Next, we applied named entity recognition (NER) models to these text chunks to identify chemical entities. In particular, we utilized an NER model [35] that leverages a PubMedBERT architecture [36] fine-tuned on various chemical datasets. To ensure that the extracted entities were specific, verified, and chemically relevant, we utilized OpenAI’s gpt-4o to review and refine the outputs. The same model was also employed to extract relations between these verified entities, forming triplets that capture the interactions and associations present in the text. Additionally, large language models were utilized to extract descriptive features from textual data associated with each entity. To enrich the nodes further during the construction of the knowledge graph, supplementary information from Wikipedia and the PubChem dataset [37] was integrated. Consequently, the finalized knowledge graph comprises nodes representing chemical entities, enhanced by metadata and descriptive annotations from these sources, as well as edges representing the relationships extracted from the textual data. Figure 1 illustrates the procedure followed to generate the knowledge graph. Refer to the Appendix for a detailed explanation of Knowledge Graph Generation.
@@ 3.2 Multi-hop Question-Answer Generation
These individual questions were then combined into a single multi-hop question using OpenAI’s o3-mini model. Importantly, the final aggregated question was constructed to begin with the last sub-question and chain the entities up to the $\text{entity}_{1}$ of the first relation, ensuring that the final answer corresponds to the answer of the first question.
During the verification phase, each one-hop question was first reviewed for clarity, relevance to chemistry, and alignment with the corresponding text that provided the answer. The multi-hop question was then assessed through an additional evaluation step, ensuring that its logical flow effectively led to the final answer. An LLM-based verification process was employed to confirm factual accuracy, answerability based on available context and metadata, and the logical coherence of the sub-questions. Feedback from domain experts was continuously incorporated into the prompts to enhance verification accuracy. To enhance the quality of the dataset, an additional round of verification was conducted on the final questions, resulting in the removal of a subset. Supplementary Section S7.3 provides more details on the types of questions that were dropped and the reasoning behind their exclusion. Figure 2 illustrates the detailed pipeline of Multi-hop QA generation. Refer to the Appendix for a detailed explanation of QA Generation.
While our pipeline is largely automated, domain expert input was integral throughout development. We conducted two rounds of pilot generations followed by detailed reviews from chemists, using their feedback to improve the pipeline and significantly enhance the quality of the generated questions. Furthermore, as explained in Section 5.1, a random subset of questions from the final dataset was evaluated by a domain expert to verify the quality of the generated dataset.
3.3 Statistical Details of the Generated Graph and Dataset
Table 1 provides a concise overview of both our dataset-level and graph-level statistics. In Table 1(a), we summarize key properties of the 971 multi-hop questions, including average question and answer lengths (in characters and tokens), the mean number of hops per question, total and pooled context lengths, and the proportion of questions containing at least one shortcut edge.
On average, questions were 319.4 chars long (SD = 129.2) or 45.5 tokens (SD = 18.6), while answers averaged just 16.7 chars (SD = 9.7) or 1.76 tokens (SD = 0.96). Each question required a mean of 2.45 hops (SD = 1.12). Summing all hops per question gives a total context length of $5,993$ chars (SD = $5,009$) or 849 tokens (SD = 726), and pooled hop lengths average $2,447$ chars (SD = $2,222$) and 346 tokens (SD = 324). Only 96 questions (9.9 %) include at least one shortcut edge (mean = 0.12, SD = 0.38). The full hop-count breakdown appears in the "Hop-count Distribution" block of Table 1(a).
QA Metric Mean Std. Dev.
Question length (chars) 319.42 129.17
Question length (tokens) 45.49 18.64
Answer length (chars) 16.66 9.69
Answer length (tokens) 1.76 0.96
Mean # hops per question 2.45 1.12
Total context length (chars) 5993.10 5009.69
Total context length (tokens) 848.55 725.78
Hop length (chars, pooled) 2447.14 2222.35
Hop length (tokens, pooled) 346.49 324.56
Shortcut count per question 0.12 0.38
Hop-count Distribution (of 971 questions)
1 hop 258 (26.6%)
2 hops 245 (25.2%)
3 hops 242 (24.9%)
4 hops 226 (23.3%)
$\geq$ 5 hops 0 (0%)
Questions w/ $\geq$ 1 shortcut 96 (9.9%)
(a) Dataset-level statistics for multi-hop questions
Graph Metric Value
Number of nodes 14 523
Number of edges 13 419
Density 0.000127
Degree (min / max / avg) 0 / 257 / 1.85
Connected components 4 684
Largest component size 7 318
Avg. clustering coefficient 0.0298
Degree assortativity coefficient –0.0265
Top 5 nodes by degree
hydrogen (257), carbon (250), oxygen (232), CO2 (220), lithium (155)
(b) Key network-level properties of the loaded knowledge graph
Table 1: Overview of both dataset-level and graph-level statistics. Left panel: Table 1(a) summarizes the dataset-level statistics for our 971 multi-hop questions. Right panel: Key properties of the underlying knowledge graph are reported in Table 1(b).
Table 1(b) then reports the main network characteristics of the underlying knowledge graph: its size (nodes and edges), sparsity (density), degree distribution (min, max, and average), number of connected components and the size of the largest component, as well as clustering and assortativity coefficients. The graph contains $14,523$ nodes and $13,419$ edges (density = 0.000127), with node degrees ranging from 0 to 257 (mean = 1.85). It splits into 4,684 connected components, the largest of which spans $7,318$ nodes. The average clustering coefficient is 0.0298, and the degree assortativity coefficient is –0.0265. Finally, the five highest-degree nodes, hydrogen, carbon, oxygen, CO2, and lithium, are listed to highlight the most central concepts in the graph.
Table 2 compares our chemical dataset (ChemKGMultiHopQA) with HotpotQA-Chemistry and ChemLitQA-multi across question count, bridged entities, entity types, answer format, domain, and source. ChemKGMultiHopQA comprises 971 ChemRxiv questions enhanced by PubChem and Wikipedia (1–4 hops, auto-built KG with expert-verified subset), offering richer multi-hop chemical reasoning than HotpotQA-Chemistry (980 Wikipedia questions, 2 hops, no chemical entities) and ChemLitQA-multi (742 ChemRxiv questions, 1 entity, LLM+expert verification).
Dataset # Qs # bridged entities # entity type Answer type Domain Source
HotpotQA-Chemistry 980 2 General Short Wikipedia Crowd (Wiki)
ChemLitQA-multi 742 1 Chemistry Long & Short ChemRxiv LLM + expert verified
ChemKGMultiHopQA 971 1–4 Chemistry Short ChemRxiv enhanced by PubChem & Wikipedia LLM + NER + KG (auto) + An expert verified subset
Table 2: Comparison of HotpotQA, ChemLitQA-multi, and ChemKGMultiHopQA data sets.
4 Experiments and Results
@@
To access the selected models for evaluation in this experiment, we used different API providers: (i) all tested OpenAI models (gpt-4o, gpt-4o-mini, o1-mini and o3-mini) are accessed via the OpenAI platform; (ii) Amazon Bedrock Platform has been used to access Anthropic Sonnet 3.7 (with and without extended thinking), Anthropic Sonnet 3.5 V2, Mistral Large, DeepSeek R1 and Llama 3.3 70B Instruct; and (iii) Google Gemma 3 27B, Qwen QwQ 32B, and DeepSeek R1 Distill Qwen 32B are accessed via the OpenRouter Platform and operate at bf16 precision. All of these models can perform function-calling tool use, so they were instructed to produce valid JSON outputs to ensure consistency and enable automated validation. All models are evaluated in two settings: with and without provided context. The first scenario reflects performance when the models are paired with an ideal retrieval-augmented generation (RAG) system, while the second scenario relies on the model’s internal memory to answer the questions. After parsing the JSON, we checked whether the output was an exact match to the ground truth. If not, OpenAI GPT-4o was instructed to perform a binary assessment, determining whether the answer was correct or not, to calculate the Correctness Rate (%) metric. Our dataset comprises 971 questions spanning 1 to 4 hops (On average, 245 questions per hop), generated using the approach described in Section 3.2. The full Q&A dataset, along with the evaluation code, is accessible here.
@@ 4.1 Models Performance
Figure 3 illustrates the performance of 13 large language models evaluated with respect to correctness rate, cost, and latency in both context-provided and context-not-provided setups. In our performance evaluation, the Llama 3.3 70B Instruct and GPT-4o models achieved the lowest cost and demonstrated notably low latency, but they also registered the lowest correctness rate, making them a cost-efficient yet less accurate options. In contrast, Claude Sonnet 3.7 (with extended thinking) achieved the highest correctness rate, albeit at the expense of significantly higher cost and latency. Meanwhile, both Qwen QWEN 32B and Deepseek R1 Distil QWEN 32B maintained a favorable balance between cost and correctness rate when context is provided – i.e., equipped with a perfect RAG system –, though they incurred above-average latency. Claude Sonnet 3.7 was found to have the highest correctness rate in none-reasoning models.
@@
Figure 3: Performance of selected models based on correctness rate, cost, and latency. The cost axis is log-scaled to emphasize differences. The y-axis represents the percentage of questions answered correctly by each model, while the size of the dots indicates the average latency for each model when answering questions. The top panel displays results for setups where context is provided to the models, while the bottom panel presents results for setups without context. Note that the horizontal axis in both panels has the same range, while the vertical axes have different ranges.
@@
In Context not provided setup, open-source reasoning models (R1 Distill Qwen, QWQ-32B, and R1) tend to have lower performance ranking compared to other models. In contrast, for OpenAI models, we observed an opposite trend, which may indicate potentially richer pre-training data. For Claude 3.7, using extended thinking did not lead to better performance compared to the no-thinking setup when the context is not provided to the model; instead, it increased the output tokens and cost. For a comprehensive breakdown of model performance and experimental settings, refer to Table 3, which details metrics such as correctness rate, latency, and token usage. Note that the output details for reasoning models also include the reasoning tokens.
4.2 Comparison with HotpotQA and ChemlitQA
@@
To demonstrate that large language models, even those designed for reasoning, often struggle with domain-specific multi-hop questions, we sampled a chemistry-related subset of HotpotQA [19], a well-known general text benchmark primarily sourced from Wikipedia. We sampled chemistry questions by starting from Wikipedia’s Chemistry category, recursively exploring its subcategories (up to three levels), and then filtering HotpotQA based on exact title matches. To maintain consistency with our evaluation scheme, we excluded distractors and included only supporting documents as context. This chemistry-specific subset of HotpotQA data is made available here.
Figure 4 illustrates the average performance of all models on each dataset under two conditions: with context provided and without context in the prompt. The results indicate that when context is provided, models achieve similar performance, with our benchmark resulting in marginally lower average performance and reduced variability among the models’ output. In the setup without context, the models found our benchmark more challenging compared to the HotpotQA chemistry subset. This observation may be due to the fact that HotpotQA was exclusively built from Wikipedia, which has been utilized in the pre-training of all evaluated models, while the new benchmark is constructed from more recent ChemRxiv papers enriched with PubChem and Wikipedia.
@@
Figure 4: Comparison of LLMs’ performance on the chemical subset of HotPotQA with the curated QA dataset from this study. Error bars reflect the standard error of the mean (S.E.M) over evaluated models.
@@ 5 Analysis and Ablation
This section presents detailed ablation studies and analyses of the benchmark. First, we will present the results of a manual evaluation conducted by our panel of domain experts on a subset of the final dataset. Next, we examine how context availability and test-time reasoning affect model performance and efficiency. Finally, we will explore how the number of reasoning hops, an indicator of question difficulty, impacts model accuracy and the number of tokens needed to generate an answer.
5.1 Expert Feedback
A panel of domain experts with PhDs in Chemistry was invited to review a randomly selected subset of questions from the database. This review focused on the accuracy and quality of both the questions and their corresponding answers, as well as the necessity of multi-hop reasoning for answering them. We started with 52 multi-hop questions, each paired with a fully worked, hop-by-hop answer. Twelve questions (23 %) were dropped due to low evaluator confidence, leaving 40 questions with high-confidence judgments. As shown in Supplementary Table S4, these 40 questions fall into three expert-rated categories: Good (26 questions, 65 %), Ok (9 questions, 22.5 %), and Poor (5 questions, 12.5 %), generally approving 87.5% of the evaluated questions. More detailed analysis is provided in section S7.4.